Solve 5/7+1/4=1/x, and what does x=?
Answer by stanbon(75887)

Solve 5/7+1/4=1/x, and what does x=?
---
lcm of 7,4,x is 28x
---
Multiply thru by 28x to get:
20x + 7x = 28
27x = 28
x = 28/27
----------
Cheers,
Stan H.